OpenStudy (anonymous):

Conduct an experiment of rolling a six-sided dice one time. Find the probability of rolling a 3 or a 4. a. 5/6 b. 1/2 c. 1/6 d. 1/3

OpenStudy (anonymous):

there are 6 possible outcomes, 2 of which satisfies the condition, therefore 2/6 = 1/3

OpenStudy (anonymous):

it would be D because you have six sides, so it means that every time you roll you can only get 1 of the 6 sides (1/6) but as you have 2 possible sides then you get 1/6+1/6=2/6 simplified into 1/3